From GlueXWiki
Jump to: navigation, search

January 31, 2013 FDC meeting


  1. Production Construction Tracking (Dave)
  2. Basis of Estimate for the FDC activities, Dave's estimate for consumables, guideline (Lubomir, Bill, Dave)
  3. Corrosion tests at MePhi Vlad's set-up, and results (Lubomir)
  4. Engineering (Bill) [1], [2], [3]
  5. Electronics (Chris, Nick)
  6. Tests at EEL126 FDC E-log (Beni)
  7. Other


Participants: Eugene, Bill, Dave, Chris, Nick, Simon, Beni, and Lubomir.


- Dave: Spare parts:about 70% are done (see the link above). Tina, Dave and Casey worked last Saturday to catch up with the foil gluing. Dave got an estimate from Vision Machine for two new spacer rings and making a grove on a type-2 cathode frame. Eugene agreed to add to this request also two spacer rings that are 3.5cm thick, these are needed for the proof-of-principle of the cluster counting with FDC. Eugene requested a page explaining the need for this.

- Lubomir for package #2: all the channels were tested and are working fine, except noise was found on wire #37 of cell #2 (same wire that was replaced before). Same noise is visible on the top and bottom strips intersecting in one point close to the middle. On Monday we opened the package and Casey tried to clean that place and we closed the package with 5 cells. On Tuesday we found the source of the noise moved by 5cm and Casey did another attempt to clean it. On Wednesday we found again it moved by 2cm and decided to replace the wire. Casey and Dave replaced and checked the wire position and tension. Interestingly, the new wire was parallel to the old ones at <10 microns level at +/-20cm from the center. That means the strong-back didn't move since this old wire frame (#10) was strung. We looked also at the wire with a microscope at that place: tiny fiber around the wire seamed to be visible and looked darker. After the meeting we installed the last two cells of the package.

Basis of Estimate for the FDC

- The BOE linked above is now more complete and includes estimates also for the installations. Bill, Dave and Lubomir discussed all the items and modified the document significantly. There are now two testing procedures for the packages. One, as it was in the original schedule "Final Assy. and Testing" which is what we have done already with two packages, i.e. partially testing using 2 fADC125. The second is the full tests with all the electronics needed for one package, that is not in the schedule. Beni will also look at the testing procedures in the BOE. Eugene requested breakdown of the man-weeks in different types (techs, scientists, engineers).

Corrosion test

- Vlad made a very good progress at MePhi with the tests. The description of his new set-up is attached. First test is at 100degC and 100% humidity and he had the two EPDM samples failed already in 6/24 hours. At Blue Crab first EPDM failed in 8 days. This could be due to the air pollution at Moscow, or that at Blue Crab we didn't have maximum humidity (Bill).

- The plan will be to continue with the test for a couple of weeks and then switch to lower temperature.


- Bill showed the pictures linked above, explaining the test of the cooling system: will have one full loop with pre-amps and will simulate the impedances of the rest. Bill is also working on the tube bending procedure.

- Survey of the installation carts at Blue Crab. Bill and Dave explained the problem: supporting the carts and different positions (where the wheels are, and 20 cm away from them) results in 0.5mm deformation of the frame at the rails. Bill is afraid that when we load the cart with all the packages and cables (155kg) it will deform and the rails will no longer be aligned. While inserting the FDC the positions of the packages as surveyed outside of the magnet, may change. As a first step, Bill wants to load a cart with 155kg and survey again the rails.


- Three pre-amp cards from the second and fourth packages have a problematic channel: either lower gain or no signal. Nick will give us new cards for replacement.

Tests at EEL126

- Beni: fixed the bad TDC and now all the wires are seen at the same time; there are two noisy wires at the present threshold (5V). Beni is taking fADC data now with 2150V/-500V and one HV channel tripped (C3HV4). The current was ~1.7 uA but went down to ~0.7uA on this channel and is stable.

- The plans is to finish with the testing of package #4 and bring package #1 to EEl126 on Monday.